If the radius of the sphere is increased by 10% keeping the mass constant then ......

Grade:12

1 Answers

Arun
6 years ago

F=ma=GMm/r²

on the surface of the earth

g=a=GM/r² , M is the mass of the earth, which is a constant in this case

g=GM/r²≈9.8 N/kg

g’=GM/(0.9r)²=100/81 g≈1.23g≈12.1 N/kg